a river has a current of 5mph. it takes al half an hour longer to paddle upstream 1.2 miles than to paddle downstream the same distance. what is Als rate in still water?
set up a quadratic equation and solve

Answer by mananth(16946)

river - 5 mph
boat x
upstream speed = x-5
downstream = x+5
1.2/(x-5)-1.2/(x+5)= 1/2
LCD = 2(x-5)(x+5)
multiply equation by LCD
2*1.2(x+5) - 2*1.2(x-5) = 1/2 * 2*(x-5)(x+5)
2.4x+12-2.4x+12=x^2-25
24+25=x^2
49=x^2
x +/- 7
speed of boat = 7 mph